For this actvity, find the matrix represenatation (T) for the linear transformation T: R3 → R2 defined by T (6) x1 + x2 -2x3help finish the matlab script
Answer #1

T: R² R² is defined as 7 90, x2 x + X2 - 223 23 T -2 Image of ordered basis of T([?]) - (****) ( [:])=(:3-6-3 T([:])= (*]n MATLAB Drive ) sample.m 1 clear all 2 clc 3% creating matrix D 4D = [3 2 1 -3 5; -15 -2 -4 0]; 5 % row reduced of augmented

Its output is

>> sample T_BtoC = 0.5294 -0.2941 -0.4118 -0.8824 1.4706 0.2941 >>

If you want output in exact form then use command

format rat

just in next line to the command clc